OpenStudy (anonymous):

Find the product of (3 + 6i)(4 - 7i)

OpenStudy (anonymous):

is it 54 + 3i?

OpenStudy (tkhunny):

Maybe. How did you get that?

OpenStudy (anonymous):

Just imputted it into my trusty calculator

OpenStudy (anonymous):

Is that correct?

OpenStudy (tkhunny):

Does your calculator lie? Can you just multiply the binomials?

OpenStudy (anonymous):

So it is lying?

OpenStudy (tkhunny):

Part of the learning of mathematics is confidence. If you perform correctly, in a manner that can be followed and reproduced, be satisfied with your result and move on. Did you do anything wrong on purpose?

OpenStudy (anonymous):

Nope it isn't. Just did the problem.

OpenStudy (anonymous):

I foiled the equation. Thanks.

OpenStudy (tkhunny):

(3 + 6i)(4 - 7i) = 12 - 21i + 24i - 42i^2 = 12 - 21i + 24i + 42 = 12 + 3i + 42 = 54 + 3i There you have it.